The mean IQ score for 2000 students is 100​, with a standard deviation of 10. Assuming the scores have a normal​ curve, answer the following.
a. How many have an IQ score between 90 and 110​?
b. How many have an IQ score between 80 and 120​?
c. How many have an IQ score over 130​?
1
Expert's answer
2020-12-15T03:38:26-0500

Given "\mu=100, \sigma=10"

a.

"P(90<X<110)=P(X<110)-P(X\leq90)"

"=P(Z<\dfrac{100-100}{10})-P(Z\leq\dfrac{90-100}{10})"

"=P(Z<0)-P(Z\leq-1)\approx0.5-0.158655"

"=0.341345"

"0.341345(2000)=683"

683 students


b.

"P(80<X<120)=P(X<120)-P(X\leq80)"

"=P(Z<\dfrac{120-100}{10})-P(Z\leq\dfrac{80-100}{10})"

"=P(Z<2)-P(Z\leq-2)\approx0.977250-0.022750"

"=0.9545"

"0.9545(2000)=1909"


By the 68-95-99.7 rule


"P(\mu-2\sigma\leq X\leq\mu+2\sigma)\approx0.9545"

"0.9545(2000)=1909"

1909 students


c.

"P(X>130)=1-P(X\leq130)"

"=1-P(Z\leq\dfrac{130-100}{10})=1-P(Z\leq 3)"

"\approx1-0.998650=0.00135"

"0.00135(2000)=3"


3 students